### v1

@@ 4 AndroidWorld as a computer-control benchmark
To test AndroidWorld applicability for autonomous agents, we develop and test a state of the art agent and its variants across all 20 apps and 116 tasks as well as on MobileMiniWoB++.
@@ 4.1.1 M3A
In the first stage, M3A generates an action, represented in JSON, and reasoning for that action. To generate this output, the agent is provided with a list of available action types, guidelines for operating the phone, and a list of UI elements derived from the Android accessibility tree’s leaf nodes. The agent receives the current screenshot and a Set-of-Mark (SoM) [62] annotated screenshot, which includes bounding boxes with numeric labels on the top-left corner for each UI element (see screenshot in Figure 5 in the Appendix). The agent attempts to execute outputted action by referencing the specific mark (if applicable). In addition to the multimodal agent, we have developed a text-only variant that consumes the screen represented using the accessibility tree and selects the relevant action in JSON format.
After executing an action, M3A reflects on its effect by observing any state changes that may have occurred. During this stage, the agent is provided with available action types, general operating guidelines, the actual action taken, and its reasoning, as well as before-and-after UI states, represented by UI element representations and screenshots with SoM annotations. We request the LLM to provide a concise summary of this step, including the intended action, success or failure, potential reasons for failure, and recommendations for subsequent actions. This summary will serve as the action history and be used for future action selection.
@@ 4.1.2 SeeAct baseline
We implement a baseline agent based on SeeAct [70], which was originally designed for GPT-4V for web navigation tasks. Specifically, we implement the best-performing variant of SeeActchoice, which grounds actions via textual choices. We implement SeeAct for the Android environment to evaluate how an existing model that performs well on web tasks [9] can be adapted and applied to Android.
To accommodate the Android environment, we adapt SeeAct in several ways. Firstly, we augment the action space from the original SeeAct implementation to support actions needed for mobile, including scroll, long press, navigate home and back, and open app actions. Secondly, in lieu of the DOM, which is not available for Android apps, we utilize the accessibility tree to construct candidate UI actions. Due to the lack of the DOM representation, we do not use the bespoke ranker model from the original implementation. However we observe that after applying a filtering heuristic to remove non-interactable elements, the majority of screens contains less than 50 candidate elements.
Table 3: Success Rates (SR) on AndroidWorld and MobileMiniWoB.
@@ 4.2 Experiment results
We evaluate the performance of M3A and SeeAct on the AndroidWorld and MobileMiniWoB++ task suites. For each task, we set the seed to 30 and provide a task-specific step budget that allows for solving the task with additional buffer steps. The agents operate in a zero-shot manner and do not draw from a memory of prior screens, although they maintain a history of prior actions and their effects. We experiment with Gemini 1.5 Pro and GPT-4 Turbo as the underlying base models. For MobileMiniWoB++, similarly to recent work, we evaluate on a subset of 62 tasks [72, 22, 15].
@@
Grounding remains a key challenge for all agents. The agents struggle with precise interactions, such as manipulating text and are often unable to recover from mistyping errors. This is evidenced by agents exhibiting better performance in data editing tasks compared to data entry tasks. As another example, operating sliders proves difficult for agents despite their simplicity for humans.
Screen understanding for mobile UIs remains challenging, with agents failing to detect subtle items that are essential for task completion (see Figure 7(a) in the Appendix). Additionally, agents struggle with certain UI patterns and affordances, often lacking the capability to sufficiently explore and adapt as humans do. For example, when encountering unfamiliar UI elements or tasks, agents are unable to figure out the correct actions through exploration (see Figure 7(c) in the Appendix). Moreover, agents sometimes struggle with tasks that simply involve confirming system states, e.g., confirming the WiFi is turned on, suggesting challenges in both task and screen understanding.
Longer tasks are harder for agents, due to increased opportunities to make mistakes. Longer tasks also may require agents to keep track of memory, and perform significant data entry, which can pose difficulties for agents. For these tasks that demand agent memory, such as performing transcriptions across apps or multiplying numbers, or tasks that involve scrolling across entries, the agents struggle as they do have little ability to “remember” what was seen on the prior screens.
Compared to M3A, SeeAct is less performant on the AndroidWorld task suite but performs similarly on MobileMiniWoB++. This stronger performance on the web is expected, as SeeAct was optimized for web environments. SeeAct struggles with operating mobile UIs, particularly with mobile-specific actions such as long-presses and swipes. Additionally, SeeAct sometimes fails to select the correct action based on its generated output, likely because it does not utilize screen elements during action generation.
Beyond difficulties with actions, SeeAct struggles with tasks requiring memory, as it only caches its actions, not their results. For instance, we observed SeeAct getting stuck in loops, issuing ineffective actions like continually scrolling despite reaching the bottom of a page. Scrolling is particularly challenging for SeeAct, as its original implementation did not require this action. Although SeeAct can sometimes recover from errors, it often does not do so quickly enough, leading to termination when the maximum number of steps is reached.
Finally, we also note the use of large foundation models introduces significant latency, and agents are about three times slower than humans to finish a task on average. On average, M3A takes 3.9 minutes to complete a task, with the text-only version taking 2.5 minutes.
@@ 4.3 Agent robustness under random seeds
Figure 3: Success rate variation across tasks. Using a fixed seed, the agent appears completely incapable of solving some tasks due to “bad luck” with the seed. In contrast under different task parameterizations, we observe the agent is capable of solving the tasks fairly often. This contrast between the two groups is reflected by higher variance for the different seed group (orange) vs. the same seed group (blue). Significant differences, with p-value <  0.05, are indicated by “*”.
We evaluate agent robustness under two conditions: (1) identical tasks with the same parameters and (2) tasks with different parameter combinations, which change the initial state and task definition. Due to computational constraints, we perform this analysis on a representative subset of AndroidWorld tasks (listed in C.4). We use the strongest agent, M3A using the accessibility tree and GPT-4, for this analysis. Our results are shown in Figure 3.
In the baseline experiment using the same seed, the agent is unable to perform the add and edit tasks, and rarely solves the two delete tasks. For the add and edit tasks, the agent struggles with UI operations, while for the delete file task, it often gets confused before the step budget is consumed. Surprisingly, the agent’s performance varies even with a fixed seed, indicating the model’s non-determinism affects agent reliability.
Under varying seeds, agent performance is much more variable, with statistically significant differences observed in the add expense and edit note tasks compared to the same seed group. Interestingly, the agent can solve the tasks fairly often under different seeds, suggesting the model’s sensitivity to task parameters and the potential for improvement via RL-like mechanisms. This experiment demonstrates the importance of supporting parameterizable tasks to better characterize real-world performance of autonomous agents.
Further, the high intra-task variation indicates current agents’ lack of robustness to environmental changes and unreliable performance due to model non-determinism. As observed in RL environments [17, 40, 8], we also observe agent performance is more accurately represented using the mean across random seeds. Finally, we note the non-zero reward observed under some seeds suggests the possibility of future improvements through RL-like mechanisms.
@@ 5 Limitations
AndroidWorld currently supports tasks with open-source Android apps with at least 1 million downloads and built-in Android system apps. While testing on more trending apps is desirable, we found open-source apps to be equally realistic and, in some cases, more challenging than apps with larger user bases. Trending apps tend to have UIs which are heavily optimized for smooth user experience, offering more UI functionality and shortcuts. Testing on less-optimized UIs makes the agent’s task harder. In an example failure we reported in the Appendix, the agent needed to delete all notes in a list and failed by repeatedly searching for a “delete-all” button. Instead, an agent with stronger reasoning capabilities would have probably searched once for that functionality, realized it was not available, and deleted the notes one by one.
@@ 6 Conclusion
We introduced AndroidWorld, a realistic and robust agent environment for Android that enables the development and evaluation of autonomous agents across a wide range of tasks and apps. AndroidWorld provides a reproducible task suite consisting of 116 tasks across 20 apps, with each task dynamically generated using random parameters to challenge agents with millions of unique goals. This dynamic nature not only tests the adaptability of agents for evaluation purposes but also opens up new research opportunities for online learning algorithms in computer control agents.
To showcase AndroidWorld’s effectiveness as a benchmark, we developed and released M3A, a state-of-the-art agent, and established benchmark results on AndroidWorld. We tested our agent using multimodal input and text-only. We found that incorporating multimodality, with Set-of-Mark prompting, does not improve performance on Android tasks and the text-only model achieved the best results. Despite outperforming a web agent adapted for Android, our best-performing agent achieved a success rate of 30.6%, which is much lower than the human success rate of 80.0%. This highlights the need for further research and development to build more reliable and robust computer control agents.
We also observe that agent performance varies significantly across variations in intent parameters, with performance even varying with the same parameters, due to unpreventable model stochasticity. These findings underscore the importance of comprehensive evaluation and the need to improve the reliability of these models under diverse real-world conditions.
AndroidWorld provides a realistic and extensible environment for developing and evaluating agents in the Android ecosystem. By releasing AndroidWorld and establishing benchmark performance with M3A, we aim to accelerate research and development in this area, ultimately leading to the creation of more reliable and robust computer control agents capable of operating effectively in real-world environments.
Appendix A AndroidWorld environment
A.1 Observation Space
@@ A.3 MobileMiniWoB++
Appendix B AndroidWorld Benchmark Details
B.1 App selection
@@
B.2 Task classification and generation
@@ 4 It checks the SMS telephony database, which is located at:
B.3 Information retrieval tasks
@@
B.4 Task examples
@@
Table 5: Examples of AndroidWorld tasks. We list the task nickname, the task template which highlight which task attributes can be parameterized, the initialization logic that is executed before the task starts and pseudo code describing the success evaluation.
Task nickname Task template Initialization logic Validation code
VlcCreatePlaylist Create a playlist in VLC, titled ”{playlist_name}” with the following files, in order: {files} Clear app state. Create new mpeg files: files + non_goal_files. Add them to VLC videos folder. execute_sql(vlc_query) == files
RecipeAddMultiple RecipesFromImage Add the recipes from recipes.jpg in Simple Gallery Pro to the recipe app. Clear app state. Write a receipt file with recipes to gallery photos. sql_rows_exist(expected_recipes)
Appendix C AndroidWorld agent details
C.1 M3A observations
@@ C.2 Agent actions
C.3 Error analysis
We analyze agent errors based on broader categories we observe during evaluation.
Grounding errors
(a) Error on initial click.
(b) Error with text entered and saving.
Figure 6: Screens illustrating grounding errors.
@@ C.3 Error analysis
C.4 Agent robustness experiments

### v2

@@ 4 AndroidWorld as a computer-control benchmark
To test AndroidWorld’s applicability for autonomous agents, we develop and test a state-of-the-art agent and its variants across all 20 apps and 116 tasks, as well as on MobileMiniWoB++.
@@ 4.1.1 M3A
In the first stage, M3A generates an action, represented in JSON, and reasoning for that action. To generate this output, the agent is provided with a list of available action types, guidelines for operating the phone, and a list of UI elements derived from the Android accessibility tree’s leaf nodes. The agent receives the current screenshot and a Set-of-Mark (SoM) [69] annotated screenshot, which includes bounding boxes with numeric labels on the top-left corner for each UI element (see screenshot in Figure 5). The agent attempts to execute outputted action by referencing the specific mark (if applicable). In addition to the multimodal agent, we have developed a text-only variant that consumes the screen represented using the accessibility tree and selects the relevant action in JSON format.
After executing an action, M3A reflects on its effect by observing any state changes that may have occurred. During this stage, the agent is provided with available action types, general operating guidelines, the actual action taken, and its reasoning, as well as before-and-after UI states, represented by UI element representations and screenshots with SoM annotations. We request the agent to provide a concise summary of this step, including the intended action, success or failure, potential reasons for failure, and recommendations for subsequent actions. This summary will serve as the action history and be used for future action selection. See App. D for more details on the agent.
@@ 4.1.2 SeeAct baseline
We implement a baseline agent based on SeeAct [78], which was originally designed for GPT-4V for web navigation. Specifically, we implement the best-performing variant, SeeActchoice, which grounds actions via textual choices. We implement SeeAct for the Android environment to evaluate how an existing model that performs well on web tasks [10] can be adapted and applied to Android.
To accommodate the Android environment, we adapt SeeAct in several ways. Firstly, we augment the action space from the original SeeAct implementation to support actions needed for mobile, including scroll, long press, navigate home and back, and open app actions. Secondly, in lieu of the DOM, which is not available for Android apps, we utilize the accessibility tree to construct candidate UI actions. Due to the lack of the DOM representation, we do not use the bespoke ranker model from the original implementation. However we observe that after applying a filtering heuristic to remove non-interactable elements, the majority of screens contains less than 50 candidate elements. See App. D.5 for more details on the implementation.
Table 3: Success Rates (SR) on AndroidWorld and MobileMiniWoB++.
@@ 4.2 Experiment results
We evaluate M3A and SeeAct on AndroidWorld and MobileMiniWoB++. We set the seed to 30 and provide a task-specific step budget. We use Gemini 1.5 Pro and GPT-4 Turbo as base models. For MobileMiniWoB++, we evaluate on a subset of 62 tasks, consistent with recent studies [80, 25, 18].
@@
4.3 Analysis
Agents have difficulty understanding mobile UIs, often failing to detect visual cues that are essential for task completion (see Figure 6(a)). Additionally, agents struggle with certain UI patterns and affordances, and when they make reasoning mistakes (see Figure 6(b)), they often lack the capability to explore and adapt as humans do (see Figure 6(c)). Moreover, agents sometimes struggle with tasks that simply involve confirming system states, e.g., confirming the WiFi is turned on, suggesting challenges in both task and screen understanding.
The agents struggle with grounding, particularly when executing precise interactions, such as manipulating text (see Figure 7) or operating sliders, and they are often unable to recover from mistyping errors. In addition, for tasks that demand memory, such as performing transcriptions across apps, multiplying numbers, or scrolling, the agents struggle as they are unable to “remember” content.
SeeAct performs less effectively than M3A on the AndroidWorld task suite and similarly on MobileMiniWoB++, reflecting its optimization for web rather than mobile environments. It struggles with mobile-specific actions like long-presses and swipes, and often fails to select appropriate actions due to not incorporating screen elements during action generation. Memory-intensive tasks are particularly challenging, as SeeAct only caches actions without remembering outcomes, leading to repetitive, ineffective behaviors such as endless scrolling. This lack of quick error recovery often results in task termination once maximum steps are reached.
Finally, we note that large foundation models significantly increase latency, taking three times longer than humans on average to complete tasks. On average, M3A takes 3.9 minutes to complete a task, with the text-only version taking 2.5 minutes.
@@ 4.4 Agent robustness under random seeds
We evaluate agent robustness under two conditions: (1) identical tasks with the same parameters and (2) tasks with different parameter combinations, which change the initial state and task definition. Due to computational constraints, we perform this analysis on a representative subset of AndroidWorld tasks (listed in D.4). We use the strongest agent, M3A using the accessibility tree and GPT-4, for this analysis. Our results are shown in Figure 3.
Figure 3: Success rate variation across tasks due to the parametrization built into AndroidWorld. Using a fixed seed, the agent appears completely incapable of solving some tasks due to “bad luck” with the seed. In contrast, under different task parameterizations, we observe the agent is capable of solving the tasks fairly often. This contrast between the two groups is reflected by higher variance for the different seed group (orange) vs. the same seed group (blue). Significant differences, with p-value <  0.05, are indicated by “*”.
In the baseline experiment with a constant seed, the agent fails the add and edit tasks, and rarely solves the two delete tasks. For the add and edit tasks, the agent struggles with UI operations, while for the delete tasks, it often gets confused before the step budget is consumed. Surprisingly, the agent’s performance varies even with a fixed seed, suggesting the model’s non-determinism affects agent reliability.
Agent performance varies much more using different seeds, with statistically significant differences for the add expense and edit note tasks. The high intra-task variation indicates the model’s sensitivity to the seed and its parameters, stemming from a lack of robustness from the underlying model.
Research in RL environments ([20, 44, 9]) has noted similar sensitivity to seeds. Consistent with these studies, we observe agent performance is best represented by the mean across random seeds. Notably, the observation of non-zero rewards under some seeds points to potential enhancements through RL-like mechanisms in future work.
Overall, this experiment underscores the importance of extensively testing agents under varied task parameters to ensure they can handle real-world variability, a capability that AndroidWorld supports effectively.
@@ 5 Limitations
AndroidWorld currently supports tasks from mostly open-source Android apps with at least 1 million downloads and built-in Android system apps. While testing on more trending apps is desirable, we found open-source apps to be equally realistic and, in some cases, more challenging than apps with larger user bases. Trending apps tend to have UIs which are heavily optimized for smooth user experience, offering more UI functionality and shortcuts. Testing on less-optimized UIs makes the agent’s task harder. In an example failure we show in Figure 6(c), the agent needed to delete all notes in a list and failed by repeatedly searching for a “delete-all” button. Instead, an agent with stronger reasoning capabilities would have probably searched once for that functionality, realized it was not available, and deleted the notes one by one.
@@ 6 Conclusion
We introduced AndroidWorld, a realistic and robust agent environment for Android that enables the development and evaluation of autonomous agents across a wide range of tasks and apps. AndroidWorld provides a reproducible task suite consisting of 116 tasks across 20 apps, with each task dynamically generated using random parameters to challenge agents with millions of unique goals. By releasing AndroidWorld and establishing benchmark performance with M3A, we aim to accelerate research and development in this area, ultimately leading to the creation of computer control agents capable of operating effectively in real-world environments. Further, the dynamic nature of AndroidWorld opens up new research opportunities for online learning algorithms in computer control agents.
Appendix A Ethical considerations
Malicious use
There is a risk that malicious actors could engineer agents to bypass security measures like CAPTCHAs or engage in activities like spamming. Additionally, they could alter prompts or screen outputs to further harmful objectives.
Societal impact
Automation agents may transform societal norms, disrupt employment, and modify human behavior. While they can enhance efficiency, this improvement could pose risks if exploited by malevolent forces.
Appendix B AndroidWorld environment
B.1 Observation Space
@@ B.3 MobileMiniWoB++
Appendix C AndroidWorld Benchmark Details
C.1 App selection
@@
C.2 Task classification and generation
@@ 4 It checks the SMS telephony database, which is located at:
26 def teardown(self, env: interface.AsyncEnv) -> None:
27 """Clears the SMS database."""
28 super().teardown(env)
29 clear_sms_database(env.base_env)
30
@@
C.3 Information retrieval tasks
@@
C.4 Humans for task analysis
During development, we recruited six software engineers from Google to analyze task difficulty, duration, and category. Each human was assigned an equal portion of tasks and tasked with identifying bugs during this annotation phase. This process resulted in the discovery and resolution of over 30 bugs.
To evaluate human performance, we enlisted two software engineers to complete the tasks using an Android emulator. Participants were provided with task descriptions and attempted to achieve the goals based on their interpretations. Each participant had one attempt per task. The majority of errors stemmed from misinterpretations or minor errors, such as entering an incorrect file extension. Other errors occurred when participants encountered unfamiliar user interfaces, impeding their ability to solve the tasks on their first attempt.
In both exercises, we informed participants about the intended use of the collected data. Participants were not required to enter any personal information in the tested tasks.
C.5 Task examples
@@
Table 5: Examples of AndroidWorld tasks. We list the task nickname, the task template indicating which task attributes can be parameterized, the initialization logic that is executed before the task starts and pseudo code describing the success evaluation.
Task nickname Task template Initialization logic Success evaluation code
VlcCreatePlaylist Create a playlist in VLC, titled “{playlist_name}” with the following files, in order: {files} Create new mpeg files: files + “noise” files that should not be added. Add them to VLC videos folder. execute_sql(vlc_query) == files
RecipeAddMultiple RecipesFromImage Add the recipes from recipes.jpg in Simple Gallery Pro to the recipe app. Write a receipt file with recipes to Simple Gallery. sql_rows_exist(expected_recipes)
MarkorEditNote Edit {file_name} in Markor. {file_operation}. Generate file with starting content, along with “noise” files not relevant to goal. Note: file_operation can be to add a footer, header, or update note content. file_exists(file_name, content=expected_content)
ExpenseAddSingle Add the following expenses into pro expense: {expense_csv} Add to the app’s SQLite database the expense that should be deleted, along with “noise” expenses that should not be deleted. sql_rows_exist(expense_obj)
SimpleCalendarDelete EventsOnRelativeDay In Simple Calendar Pro, delete all events scheduled for this {day_of_week}. add to the app’s SQLite database calendar events on specified day, along with “noise” events that should not be deleted. !sql_rows_exist(expected_events)
FilesDeleteFile Delete the file {file_name} from the Android filesystem located in the {subfolder} folder within the sdk_gphone_x86_64 storage area. Generate specified file, along with “noise” files that should not be deleted. !file_exists(file_name)
SportsTrackerActivities CountForWeek How many {category} activities did I do this week in the OpenTracks app? Express your answer as a single integer. add to the app’s SQLite database activities for the specified category, along with “noise” activities. int(agent_response) == expected_count
Appendix D AndroidWorld agent details
D.1 M3A observations
@@ D.2 M3A actions
D.3 Error analysis
We analyze M3A errors based on broader categories we observe during evaluation.
@@ D.3 Error analysis
Grounding errors
(a) Error on initial click.
(b) Error with text entered and saving.
Figure 7: Screens illustrating grounding errors.
D.4 Agent robustness experiments
@@
More details about these tasks can be found in Table 5.
D.5 SeeAct details
We modify the SeeAct prompt [78] to reflect that the environment is Android by inputting elements from the accessibility tree and supporting additional actions (e.g., scrolling). Below we include the updated prompt. We annotate the system, user, and assistant roles that are each provided to the OpenAI API.
1
2> Role: SYSTEM
3Imagine that you are imitating humans operating an Android device for a task step by step. At each stage, you can see the Android screen like humans by a screenshot and know the previous actions before the current step decided by yourself through recorded history. You need to decide on the first following action to take. You can tap on an element, long-press an element, swipe, input text, open an app, or use the keyboard enter, home, or back key. (For your understanding, they are like adb shell input tap’, adb shell input swipe’, adb shell input text’, adb shell am start -n’, and adb shell input keyevent’). One next step means one operation within these actions. Unlike humans, for typing (e.g., in text areas, text boxes), you should try directly typing the input or selecting the choice, bypassing the need for an initial click. You should not attempt to create accounts, log in or do the final submission. Terminate when you deem the task complete or if it requires potentially harmful actions.
4
5> Role: USER
6You are asked to complete the following task: <GOAL>
7
8Previous Actions:
9<PREVIOUS ACTIONS>
10
11The screenshot below shows the Android screen you see. Follow the following guidance to think step by step before outlining the next action step at the current stage:
12
13(Current Screen Identification)
14Firstly, think about what the current screen is.
15
16(Previous Action Analysis)
17Secondly, combined with the screenshot, analyze each step of the previous action history and their intention one by one. Particularly, pay more attention to the last step, which may be more related to what you should do now as the next step. Specifically, if the last action involved a INPUT TEXT, always evaluate whether it necessitates a confirmation step, because typically a single INPUT TEXT action does not make effect. (often, simply pressing Enter’, assuming the default element involved in the last action, unless other clear elements are present for operation).
18
19(Screenshot Details Analysis)
20Closely examine the screenshot to check the status of every part of the screen to understand what you can operate with and what has been set or completed. You should closely examine the screenshot details to see what steps have been completed by previous actions even though you are given the textual previous actions. Because the textual history may not clearly and sufficiently record some effects of previous actions, you should closely evaluate the status of every part of the screen to understand what you have done.
21
22(Next Action Based on Android screen and Analysis)
23Then, based on your analysis, in conjunction with human phone operation habits and the logic of app design, decide on the following action. And clearly outline which element on the Android screen users will operate with as the first next target element, its detailed location, and the corresponding operation.
24
25To be successful, it is important to follow the following rules:
261. You should only issue a valid action given the current observation.
272. You should only issue one action at a time
283. For handling the select dropdown elements on a screen, its not necessary for you to provide completely accurate options right now. The full list of options for these elements will be supplied later.
29
30> Role: ASSISTANT
31<AGENT RESPONSE TO ABOVE>
32
33> Role: USER
34(Reiteration)
35First, reiterate your next target element, its detailed location, and the corresponding operation.
36
37(Multichoice Question)
38Below is a multi-choice question, where the choices are elements on the screen. All elements are arranged in the order based on their height on the screen, from top to bottom (and from left to right). This arrangement can be used to locate them. From the screenshot, find out where and what each one is on the screen, taking into account both their text content and details. Then, determine whether one matches your target element. Please examine the choices one by one. Choose the matching one. If multiple options match your answer, choose the most likely one by re-examining the screenshot, the choices, and your further reasoning. If you would like to perform a swipe action, you can optionally select the choice where you will swipe.
39
40A. "Home" icon
41B. "Phone" icon
42C. "Messages" icon
43D. "Chrome" icon
44E. "Search" icon
45...
46If none of these elements match your target element, please select Z. None of the other options match the correct element.
47
48(Final Answer)
49Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element choice, action, and value should be in three separate lines.
50
51Format:
52
53ELEMENT: The uppercase letter of your choice. (No need for TERMINATE, KEYBOARD ENTER, WAIT, ANSWER, OPEN APP, NAVIGATE HOME, NAVIGATE BACK; and optional for SWIPE.)
54
55ACTION: Choose an action from {CLICK, INPUT TEXT, LONG PRESS, NAVIGATE BACK, TERMINATE, KEYBOARD ENTER, SWIPE, WAIT, ANSWER, OPEN APP, NAVIGATE HOME}.
56
57VALUE: Provide additional input based on ACTION.
58
59The VALUE means:
60If ACTION == INPUT TEXT, specify the text to be typed.
61If ACTION == SWIPE, specify the direction: up, down, left, right.
62If ACTION == OPEN APP, provide the name of the app to be opened.
63If ACTION == ANSWER, specify the text of your answer to respond directly to a question or request for information.
64For CLICK, LONG PRESS, KEYBOARD ENTER, NAVIGATE HOME, NAVIGATE BACK, WAIT, and TERMINATE, write "None".